OpenStudy (anonymous):

Please help me!! Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

is it \(\cot(x)\sec^4(x)\)?

OpenStudy (anonymous):

@satellite73 yes!!!!

OpenStudy (anonymous):

ok then the left hand side can be rewritten as \[\cot(x)\sec^4(x)=\csc(x)\sec^3(x)\] since \(\cot(x)=\frac{\cos(x)}{\sin(x)}\) as one of the cosines cancel now lets see what we can do to the right hand side

OpenStudy (anonymous):

@satellite73 okay so the right side will be {cos(x)/sin(x)}+2tanx+tan^3(x)?? btw does that sat sec to the second power or third power?

OpenStudy (anonymous):

this is a raft of algebra

OpenStudy (anonymous):

let me see if i can straighten in out clearly before i start writing

OpenStudy (anonymous):

the idea will be to replace \[2\tan(x)+\tan^3(x)=\tan(x)(2+\tan^2(x))=\tan(x)(1+1+\tan^2(x))\] \[=\tan(x)(1+\sec^2(x))\] and then add the fractions

OpenStudy (anonymous):

@satellite73 could you work out the whole solution so i could get an idea of it?O.o

OpenStudy (anonymous):

yes i will as soon as i get it you are sure there is not a typo here right? because i am getting very close, but not getting it exactly

OpenStudy (anonymous):

@satellite73 hmm what typo could have been made?cot x sec^4(x) = cot x + 2 tan x + tan^3(x) does this look better??:)

OpenStudy (anonymous):

ok i am not getting it exactly in fact i am not sure it is right lets do some algebra

OpenStudy (anonymous):

@satellite73 okay!!

OpenStudy (anonymous):

it is definitely \[\cot(x)\sec^4(x)=\cot(x)+2\tan(x)+\tan^3(x)\] right?

OpenStudy (anonymous):

@satellite73 yupp

OpenStudy (anonymous):

ok lets replace \(\cos(x)\) by \(a\) and \(\sin(x)\) by \(b\) to make the algebra simpler then if we see \(a^2+b^2\) we can replace it by 1 since \(\cos^2(x)+\sin^2(x)=1\)

OpenStudy (anonymous):

the left side is \[\frac{a}{b}\times \frac{1}{a^4}=\frac{1}{a^3b}\]

OpenStudy (anonymous):

@satellite73 yea

OpenStudy (anonymous):

ok now before we start on the right hand side, lets note that \[2\tan(x)+\tan^3(x)=\tan(x)\left(2+\tan^2(x)\right)=\tan(x)\left (1+1+\tan^2(x)\right)\] \[=\tan(x)\left(1+\sec^2(x)\right)\] since \(1+\tan^2(x)=\sec^2(x)\)

OpenStudy (anonymous):

now returning to \(a\) and \(b\) the right hand side is \[\frac{a}{b}+\frac{b}{a}\left(1+\frac{1}{a^2}\right)\] with me so far?

OpenStudy (anonymous):

@satellite73 yes!!!:D

OpenStudy (anonymous):

now we multiply out and add the fractions

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{b}{a}\left(1+\frac{1}{a^2}\right)\] \[=\frac{a}{b}+\frac{b}{a}+\frac{b}{a^3}\] adding the first two fractions gives \[\frac{a^2+b^2}{ab}+\frac{b}{a^3}\]

OpenStudy (anonymous):

now w recall that \(a^2+b^2=\cos^2(x)+\sin^2(x)=1\) so we have \[\frac{1}{ab}+\frac{b}{a^3}\]

OpenStudy (anonymous):

adding this fractions, the common denominator is \(a^3b\) so we get \[\frac{a^2}{a^3b}=\frac{b^2}{a^3b}=\frac{a^2+b^2}{a^3b}\]

OpenStudy (anonymous):

typo there sorry, it is plus, not equal \[\frac{a^2}{a^3b}+\frac{b^2}{a^3b}=\frac{a^2+b^2}{a^3b}\]

OpenStudy (anonymous):

now once again we remember that \(a^2+b^2=1\) giving us \[\frac{1}{a^3b}\] which is exactly what we wanted

OpenStudy (anonymous):

@satellite73 so what would be the final answer?

OpenStudy (anonymous):

this was a long one, but most of it was just algebra

OpenStudy (anonymous):

well you had to show that both sides were the same, and that is what we did

OpenStudy (anonymous):

the right hand side is \[\frac{1}{\cos^3(x)\sin(x)}\] and so is the left hand side

OpenStudy (anonymous):

i could try to write this out with \(\cos(x)\) and \(\sin(x)\) but it would take forever. there were two tricks here the first was to call \(2+\tan^2(x)=1+1+\tan^2(x)=1+\sec^2(x)\) and the rest was algebra, but recalling that \(a^2+b^2=1\)

OpenStudy (anonymous):

i have all the steps written out without leaving any thing out copy it down exactly, but replace \(a\) by \(\cos(x)\) and \(b\) by \(\sin(x)\) and you will have all the steps

OpenStudy (anonymous):

for example \[\frac{a}{b}+\frac{b}{a}\left(1+\frac{1}{a^2}\right)\] will become \[\frac{\cos(x)}{\sin(x)}+\frac{\sin(x)}{\cos(x)}\left(1+\frac{1}{\cos^2(x)}\right)\] and so on

OpenStudy (anonymous):

the "final answer" is really the worked out algebra, because your job was to show that the right and left hand sides were the same thing